# TDD Without Tears: Towards Test Case Generation from Requirements through Deep Reinforcement Learning

### v1

@@ Abstract.
Test-driven development (TDD) is a widely-employed software development practice that mandates writing test cases based on requirements before writing the actual code. While writing test cases is the centerpiece of TDD, it is time-consuming, expensive, and often shunned by developers. To address these issues associated with TDD, automated test case generation approaches have recently been investigated. Such approaches take source code as input, but not the requirements. Therefore, existing work does not fully support true TDD, as actual code is required to generate test cases. In addition, current deep learning-based test case generation approaches are trained with one learning objective, i.e., to generate test cases that are exactly matched with the ground-truth test cases. However, such approaches may limit the model’s ability to generate different yet correct test cases. In this paper, we introduce PyTester, a Text-to-Testcase generation approach that can automatically generate syntactically correct, executable, complete, and effective test cases while being aligned with a given natural language requirement. We evaluate PyTester on the public APPS benchmark dataset, and the results show that our Deep RL approach enables PyTester, a small language model, to outperform much larger language models like GPT3.5, StarCoder, and InCoder. Our findings suggest that future research could consider improving small over large LMs for better resource efficiency by integrating the SE domain knowledge into the design of reinforcement learning architecture.
booktitle: Proceedings of the 32nd ACM Symposium on the Foundations of Software Engineering (FSE’24), November 15–19, 2024, Porto de Galinhas, Brazilcopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Software and its engineering Software development techniques
1. Introduction
@@
TDD is an agile software development practice (Beck et al., 2001) that advocates the creation of test cases before code development (Beck, 2022; Astels, 2003). The methodology comprises iterative steps of creating test cases, writing code to pass the test cases, and refactoring. Thus, developers can conceptualize the specification before coding, which increases developers’ productivity (Erdogmus et al., 2005). Therefore, writing high-quality software test cases is crucial for TDD. However, writing test cases is a time-consuming part of TDD. To address this issue, previous studies proposed various automated test case generation approaches to automatically generate test cases based on random (Pacheco et al., 2007; Pacheco and Ernst, 2007), search algorithms (Lukasczyk and Fraser, 2022; Fraser and Arcuri, 2011), and deep learning (Watson et al., 2020; Tufano et al., 2020; Alagarsamy et al., 2023) techniques.
Nonetheless, there are two main limitations to the previous approaches. First, existing test case generation approaches often take code as input, which is not applicable to TDD, as TDD mandates writing test cases before writing code. Second, existing test case generation approaches often aim to generate test cases that exactly match the ground truth test cases, without considering alternative test cases, their characteristics (i.e., syntax correctness, executability, and completeness) of test cases, and importantly, the alignment of the test cases to the given requirement.
@@
In this paper, we propose PyTester, a Text-to-Testcase generation approach for Python using Deep Reinforcement Learning. Our Deep RL-based model is optimized with Proximal Policy Optimization (PPO) on a reward function that considers three types of feedback, namely, syntax correctness, test executability, and code coverage. We evaluate our approach on four evaluation aspects, i.e., syntax correctness, test completeness measured by code coverage, test effectiveness measured by mutation score, and requirement alignment. Evaluating the alignment of the test cases to a given requirement is indeed challenging. To address this challenge, we define that a test case is aligned with a given requirement if the test case is executable (i.e., all assertions are passed.) with the ground truth code, assuming that the ground truth code is the correct implementation of the given requirement. Then, we compare our PyTester with four strong large language models (LLMs), namely, CodeT5-large(Le et al., 2022), GPT3.5(OpenAI, 2023), StarCoder (Li et al., 2023), and InCoder (Fried et al., 2022). Through a comprehensive experiment on the APPS benchmark dataset (Hendrycks et al., 2021), we address the following research questions:
  1. (RQ1)
@@
    Of the test cases generated by PyTester, 99% are syntactically correct, 84% are aligned with the requirements while achieving a code coverage of 80% and a mutation score of 61%, which outperforms all of the studied LLMs. This finding demonstrates that a smaller language model like PyTester can outperform larger language models when carefully designed. Importantly, after being optimized with deep reinforcement learning, PyTester outperforms GPT3.5 for all evaluation metrics, highlighting the effectiveness of our deep reinforcement learning framework for Text-to-Testcase generation.
  2. (RQ2)
@@
    The top-3 reward functions for our PyTester are the Syntax+Coverage, Syntax+Executability, and Syntax-only types of feedback. Without considering syntax correctness, model performance decreases by 3-13%, indicating that syntax correctness must be the minimum consideration when designing a reward function for deep reinforcement learning in the Text-to-Testcase generation task.
  3. (RQ3)
@@
These results lead us to conclude that the performance of smaller language models like PyTester is not necessarily inferior to that of large language models if carefully designed. Our results confirms that Deep Reinforcement Learning substantially contributes to performance improvements. Deep RL enables smaller language models, with the smallest model parameter size and the fastest inference times by at least one order of magnitude, to outperform much larger language models. Furthermore, our findings indicate that incorporating test case characteristics into the reward function significantly enhances performance, resulting in the model generating test cases that are 44% more aligned with the requirement with 44% higher code coverage and 32% higher mutation score when compared to the GPT3.5 large language model. Finally, these findings emphasize the importance of considering domain knowledge and test case characteristics when designing an RL-based Text-to-Testcase generation approach.
@@
  • We, conceptually, introduce PyTester, a Text-to-Testcase generation approach that can automatically generate syntactically correct, executable, complete, and effective test cases while being aligned with a given natural language requirement.
  • We, technically, formulate the Text-to-Testcase generation task as a deep reinforcement learning problem where the reward function is designed to consider various characteristics of test cases, including syntax correctness, test executability, and code coverage.
  • We, empirically, demonstrate that our PyTester model outperforms all of the studied state-of-the-art Large Language Models including GPT3.5, StarCoder, and InCoder, while being the smallest language model with the quickest inference time.
To support the open science initiatives and increase the verifiability of our study, the replication package of PyTester is available at https://github.com/tddpytester/pytester.
Paper Organization. The paper is organized as follows. Section 2 describes the background and motivation of our study. Section 3 presents our PyTester approach. Section 4 presents the experimental setup and the model experimentation. Section 5 discusses the results to answer our research questions. Section 6 discusses related work about reinforcement learning in software engineering. Section 7 discloses the threats to validity. Section 8 draws the conclusion.
2. Background and Motivation
In this section, we describe the background of TDD, discuss existing automated test case generation approaches, and a motivating example of using ChatGPT for Text-to-Testcase generation.
2.1. Test-Driven Development (TDD)
Test-Driven Development (TDD) (Beck, 2022; Astels, 2003) is a software development methodology that involves writing tests for a piece of code before actually implementing the code itself. Aligned with the Shift-Left Testing principles and Agile manifesto (Beck et al., 2001; Ambler, 2004), TDD is typically carried out in small, incremental steps, with the developer 1⃝ writing a failing test based on requirements, then 2⃝ writing the minimum amount of code required to make the test pass, and finally 3⃝ refactoring the code as needed. With TDD, developers need to first understand the requirements and think about the expected behavior of their code, leading to more reliable and maintainable code. Prior studies also found numerous benefits of TDD, including increasing requirements understanding (Mueller and Hagner, 2002), producing more clean code with a higher number of test cases (Janzen and Saiedian, 2008; Erdogmus et al., 2005). However, the process of writing test cases is still manual, time-consuming, and labor-intensive. This limitation highlights the need for an automated test case generation approach based on a given requirement, which likely speeds up the TDD process and improves developers’ productivity.
2.2. Automated Test Case Generation Approaches
@@
Limitations. While different test case generation approaches are proposed, none of these approaches are designed for TDD (i.e., taking requirement as input), are not able to generate test cases without code (i.e., mostly they take code as input), are mostly designed for Java—not Python, the dynamic type programming language. Therefore, these limitations highlight the need for an automated test case generation approach that takes a requirement as input and can generate alternative test cases beyond the ones that exactly match the ground truth test cases.
2.3. ChatGPT for Text-to-Testcase Generation: Motivation
@@
ChatGPT (OpenAI, 2023) is a powerful AI conversational large language model (LLM) designed for generating natural text based on a given prompt question in a dialogue style. Recently, researchers found that ChatGPT can perform very well in various software engineering tasks (Hou et al., 2023), including code generation (Chen et al., 2022) and test case generation (Schäfer et al., [n. d.]; Siddiq et al., 2023; Tang et al., 2023; Xie et al., 2023; Yuan et al., 2023). For example, Chen et al. (2022) proposed CodeT, a code generation approach based on a heuristic ranking process using a dual execution agreement of ChatGPT-generated codes and test cases. Schäfer et al. ([n. d.]) proposed TestPilot, a ChatGPT-based test case generation approach for JavaScript based on a given input function. Siddiq et al. (2023) investigated the performance of LLMs for test case generation based on a Java focal method (i.e., methods to be tested). Similarly, many recent work (Tang et al., 2023; Xie et al., 2023; Yuan et al., 2023) also investigated the performance of ChatGPT-based test case generation for a given Java focal method.
@@
While ChatGPT has been widely used for test case generation, existing work only focuses on taking code as input—not the requirements as input. Therefore, existing test case generation approaches do not support TDD, as TDD mandates writing test cases based on requirements before writing the actual code. In addition, a systematic literature review by Hou et al. (2023) also confirmed that test case generation approaches to support TDD (i.e., LLM-based Text-to-Testcase generation) remain largely unexplored. To address this research gap, we first conduct a preliminary analysis to empirically investigate the performance of ChatGPT for Text-to-Testcase generation.
@@
Figure 1. An example of test cases generated by the OpenAI’s GPT3.5 model based on a given requirement input. Syntax correctness is measured by an AST parser, while the test executability is measured by running the test cases against the ground truth code.
@@
An Illustrative Example. We use ChatGPT through the OpenAI API by selecting the recommended GPT-3.5-turbo model endpoint. [^1] The experiment was run in June 2023. Following the prompt engineering’s best practices [^2] https://www.promptingguide.ai/ and Chen et al. (2022), we design a prompt that includes the context (i.e., external information or additional context that can steer the model to better responses) and the instruction (i.e., a specific task or instruction that the model should perform). Figure 1 (left) presents an example of the input format (i.e., prompt format) where the context refers to the signature of the Python function with a natural language requirement in the docstring, without the code implementation (Figure 1 (top right)), while the instruction refers to a natural language comment “# check the correctness of the ‘…’ function” followed by an incompleted assert statement. We randomly select one programming task (Problem #4479) from the testing set of the APPS dataset (Hendrycks et al., 2021) as an illustrative example. Note that a comprehensive evaluation of ChatGPT is provided in Section 5 where we found that our PyTester performs better. For the given Problem #4479, we reformat the data structure into a triplet of $<$function signature + requirement text, code implementation, test cases$>$. Then, the function signature + requirement text is input into ChatGPT to generate test cases, while code implementation is only used to evaluate if the generated test cases can be run against the ground truth code.
@@
Observation-1: ChatGPT can generate syntactically correct test cases, yet some of the generated test cases cannot be executed (i.e., failed test cases). In reality, when using ChatGPT to generate test cases based on a given requirement, it becomes challenging to verify if the generated test cases are correct and aligned with the given requirement or not. To address this challenge, assuming that the ground truth code is the correct implementation, we execute the test cases against the ground truth code to verify if the ChatGPT-generated test cases are executable or not. Figure 1 (bottom right) shows that four test cases generated by ChatGPT are syntactically correct, but the fourth test case is not executable. In particular, the fourth test case is generated with the test input of ‘([-2,9,-8,8,7], 3)’. Thus, the expected output should be ‘30’, but incorrectly generated as ‘32’. Therefore, this test case will raise an AssertionError. Such syntactically correct, yet non-executable test cases indicate that the test case is still not aligned with the given requirements. This observation demonstrates that ChatGPT is able to generate syntactically correct test cases, yet fails to generate executable ones (i.e., not aligned with the requirements).
@@
Observation-2: When ChatGPT generates non-executable test cases, it becomes challenging to evaluate the completeness and effectiveness of the test cases. Intuitively, effective test cases should exercise all lines of code in the program (i.e., achieving high code coverage), and be able to discover bugs/defects when the program is slightly changed (i.e., the same set of test cases is able to catch the introduced bugs). Code coverage is widely used to measure test completeness  (Pacheco and Ernst, 2007; Fraser and Arcuri, 2011; Schäfer et al., [n. d.]; Siddiq et al., 2023; Tufano et al., 2020; Alagarsamy et al., 2023; Tang et al., 2023). Thus, a higher code coverage percentage often increases the confidence that the software behaves as intended. However, studies (Hemmati, 2015; Cai and Lyu, 2005) argue that code coverage only measures the extent to which the test cases exercise the lines of code in the program, but not the effectiveness of the test cases. Therefore, mutation testing is proposed to evaluate the effectiveness of the test cases (Cai and Lyu, 2005; Madeyski, 2010; Jia and Harman, 2010; Papadakis et al., 2019). Mutation testing aims to introduce intentional faults (mutations) into the code and check if the existing test cases can detect these faults (mutants)—referring to the fault detection capability of the test cases. Thus, the mutation score is measured as a percentage based on the number of mutants that are killed (i.e. detected by the test cases) versus the total number of mutants introduced. Despite the importance and benefits of code coverage and the extensive research on mutation testing in the literature, we found that, both code coverage and mutation score have not been taken into consideration by existing techniques on ChatGPT-based test case generation.
@@
Summary. TDD offers numerous benefits to software development, yet is time-consuming and expensive. Existing automated test case generation approaches do not yet support TDD practices (i.e., code, not requirements, is treated as input). Our motivating analysis also shows that ChatGPT (the most powerful language model) still generates non-executable, incomplete, and ineffective test cases, hindering the adoption of ChatGPT and other LLMs for Text-to-Testcase generation in practice. To support TDD practices, this observation highlights the urgent need for a novel automated test case generation approach that is able to generate syntactically correct, executable, complete, and effective test cases.
3. PyTester

### v2

# PyTester: Deep Reinforcement Learning for Text-to-Testcase Generation

@@ Abstract
Test-driven development (TDD) is a widely-employed software development practice that mandates writing test cases based on a textual description before writing the actual code. While writing test cases is the centerpiece of TDD, it is time-consuming, expensive, and often shunned by developers. To address these issues associated with TDD, automated test case generation approaches have recently been investigated. Such approaches take source code as input, but not the textual description. Therefore, existing work does not fully support true TDD, as actual code is required to generate test cases. In addition, current deep learning-based test case generation approaches are trained with one learning objective, i.e., to generate test cases that are exactly matched with the ground-truth test cases. However, such approaches may limit the model’s ability to generate different yet correct test cases. In this paper, we introduce PyTester, a Text-to-Testcase generation approach that can automatically generate syntactically correct, executable, complete, and effective test cases while being aligned with a given textual description. We evaluate PyTester on the public APPS benchmark dataset, and the results show that our Deep RL approach enables PyTester, a small language model, to outperform much larger language models like GPT3.5, StarCoder, and InCoder. Our findings suggest that future research could consider improving small over large LMs for better resource efficiency by integrating the SE domain knowledge into the design of reinforcement learning architecture.
keywords:
Text-To-Testcase Generation , Deep Reinforcement Learning
journal: Journal of Systems and Software (JSS)
[…]
1 Introduction
@@
TDD is an agile software development practice (Beck et al., 2001) that advocates the creation of test cases before code development (Beck, 2022; Astels, 2003). The methodology comprises iterative steps of creating test cases, writing code to pass the test cases, and refactoring. Thus, developers can conceptualize the specification before coding, which increases developers’ productivity (Erdogmus et al., 2005). Therefore, writing high-quality test cases is crucial for TDD, yet still time-consuming and complex.
Following TDD practices, developers may focus on writing test cases at different granularity levels (e.g., acceptance, integration, system, and unit test cases) depending on the details of the requirements (e.g., user stories, system description, class description, function description). However, current research only supports a few scenarios of TDD (e.g., generating acceptance test cases from natural language specifications). For example, Fischbach et al. (2023) and Wang et al. (2020) proposed an approach to generate acceptance test cases based on a use case specification. However, these approaches only generate test cases at the acceptance level for a given use case. While generating acceptance test cases will greatly help teams to check if all of the use cases are met, it does not guarantee functionality correctness for every method in the software.
Therefore, researchers proposed various automated test case generation approaches to automatically generate test cases at the unit testing level using various techniques, e.g., random techniques (Pacheco et al., 2007; Pacheco and Ernst, 2007), search-based techniques (Lukasczyk and Fraser, 2022; Fraser and Arcuri, 2011), and deep learning techniques (Watson et al., 2020; Tufano et al., 2020; Alagarsamy et al., 2023). However, these test case generation approaches often take code as input, which does not apply to TDD, as TDD mandates writing test cases before coding. Thus, it remains challenging for developers to follow the TDD practices, i.e., writing unit tests based on a given textual description (i.e., the function description). Therefore, a novel Text-to-Testcase generation approach is required to help developers write unit tests based on a given textual description, which may improve the overall efficiency of automated software testing.
@@
In this paper, we propose PyTester, a Text-to-Testcase generation approach for Python using Deep Reinforcement Learning. Our Deep RL-based model is optimized with Proximal Policy Optimization (PPO) on a reward function that considers three types of feedback, namely, syntax correctness, test executability, and code coverage. We evaluate our approach on four evaluation aspects, i.e., syntax correctness, passing test case, test completeness measured by code coverage, and test effectiveness measured by mutation score. Evaluating the alignment of the test cases to a given textual description is indeed challenging. To address this challenge, we define that a test case is aligned with a given textual description if the test case is executable (i.e., all assertions are passed.) with the ground truth code, assuming that the ground truth code is the correct implementation of the given textual description. Then, we compare our PyTester with five strong large language models (LLMs), namely, CodeT5-large(Le et al., 2022), GPT3.5(OpenAI, 2023), Copilot(GitHub, 2021), StarCoder (Li et al., 2023), and InCoder (Fried et al., 2022). Through a comprehensive experiment on the APPS benchmark dataset (Hendrycks et al., 2021), we address the following research questions:
  1. RQ1.
@@
    Of the test cases generated by PyTester, 99% are syntactically correct, 84% are passed (i.e., correct test cases) while achieving a code coverage of 80% and a mutation score of 61%, which outperforms all of the studied LLMs. This finding demonstrates that a smaller language model like PyTester  can outperform larger language models when carefully designed. Importantly, after being optimized with Deep Reinforcement Learning, PyTester outperforms GPT3.5 for all evaluation metrics, highlighting the effectiveness of our Deep RL framework for Text-to-Testcase generation.
  2. RQ2.
@@
    The top-3 reward functions for our PyTester are the Syntax+Coverage, Syntax+Executability, and Syntax-only types of feedback. Without considering syntax correctness, model performance decreases by 3-13%, indicating that syntax correctness must be the minimum consideration when designing a reward function for Deep RL in the Text-to-Testcase generation task.
  3. RQ3.
@@
These results lead us to conclude that the performance of smaller language models like PyTester is not necessarily inferior to that of large language models if carefully designed. Our results confirm that Deep Reinforcement Learning substantially contributes to performance improvements. Deep RL enables smaller language models, with the smallest model parameter size and the fastest inference times by at least one order of magnitude, to outperform much larger language models. Furthermore, our findings indicate that incorporating test case characteristics into the reward function significantly enhances performance, resulting in the model generating test cases that are 6% more syntactically correct, 5% more aligned with the description (i.e., passing test case) with 6% higher code coverage and 7% higher mutation score when compared to the GPT3.5 large language model. These findings emphasize the importance of considering domain knowledge and test case characteristics when designing a Deep RL-based Text-to-Testcase generation approach.
Finally, we acknowledge the importance of the requirement understanding to generate test cases by software engineers. Thus, the main purpose of PyTester is only to assist the TDD process during the test case generation step—neither to fully automate the TDD process nor replace human involvement. Although our PyTester can generate test cases based on a given textual description, it remains essential for software engineers to review and refine these generated test cases to ensure their correctness and completeness.
@@
  • margin:
  • We, conceptually, introduce PyTester, a Text-to-Testcase generation approach for a single call function that can automatically generate syntactically correct, executable, complete, and effective test cases while being aligned with a given textual description.
  • We, technically, formulate the Text-to-Testcase generation task as a deep reinforcement learning problem where the reward function is designed to consider various characteristics of test cases, including syntax correctness, test executability, and code coverage. margin:
  • We, empirically, demonstrate that on the APPS dataset, our PyTester model outperforms all of the studied state-of-the-art Large Language Models including GPT3.5, Copilot, StarCoder, and InCoder, while being the smallest language model with the fastest inference time.
To support the open science initiatives and increase the verifiability of our study, the replication package of PyTester is available at https://github.com/awsm-research/pytester.
2 Background and Motivation
In this section, we describe the background of TDD, discuss existing automated test case generation approaches and a motivating example of ChatGPT for Text-to-Testcase generation.
2.1 Test-Driven Development (TDD)
Test-Driven Development (TDD) (Beck, 2022; Astels, 2003) is a software development methodology that involves writing tests for a piece of code before actually implementing the code itself. Aligned with the Shift-Left Testing principles and Agile manifesto (Beck et al., 2001; Ambler, 2004), TDD is typically carried out in small, incremental steps, with the developer ① writing a failing test based on textual descriptions, then ② writing the minimum amount of code required to make the test pass, and finally ③ refactoring the code as needed. With TDD, developers need to first understand the descriptions and think about the expected behavior of their code, leading to more reliable and maintainable code. Prior studies also found numerous benefits of TDD, including increasing requirements understanding (Mueller and Hagner, 2002), and producing more clean code with a higher number of test cases (Janzen and Saiedian, 2008; Erdogmus et al., 2005). However, the process of writing test cases remains manual, time-consuming, and labor-intensive. This limitation highlights the need for an automated test case generation approach based on textual descriptions, which could potentially accelerate the TDD process and enhance developer productivity.
2.2 Automated Test Case Generation Approaches
@@
margin:
NLP-based test case generation (e.g., Jdoctor (Blasi et al., 2018), and TOGA (Dinella et al., 2022)) applied pattern, lexical, and semantic matching to translate Javadoc comments into test cases (Jdoctor) or a unified transformer-based approach to translate the context of a Java Method into assertion tests (TOGA). However, these approaches are limited to Java programming language.
@@
Limitations. While different test case generation approaches are proposed, none of these approaches are designed for TDD (i.e., taking textual description as input), are not able to generate test cases without code (i.e., mostly they take code as input) and are mostly designed for Java—not Python. Therefore, these limitations highlight the need for a Python automated test case generation approach that takes a description as input and can generate alternative test cases beyond the ones that exactly match the ground truth test cases.
2.3 ChatGPT for Text-to-Testcase Generation: Motivation
@@
ChatGPT (OpenAI, 2023) is a powerful AI conversational large language model (LLM) designed for generating natural text based on a given prompt question in a dialogue style. Recently, researchers found that ChatGPT can perform very well in various software engineering tasks (Hou et al., 2023), including code generation (Chen et al., 2022) and test case generation (Schäfer et al., ; Siddiq et al., 2023; Tang et al., 2023; Xie et al., 2023; Yuan et al., 2023). For example, Chen et al. (2022) proposed CodeT, a code generation approach based on a heuristic ranking process using a dual execution agreement of ChatGPT-generated codes and test cases. Schäfer et al. proposed TestPilot, a ChatGPT-based test case generation approach for JavaScript based on a given input function. Siddiq et al. (2023) investigated the performance of LLMs for test case generation based on a given Java focal method (i.e., methods to be tested). Similarly, many recent work (Tang et al., 2023; Xie et al., 2023; Yuan et al., 2023) also investigated the performance of ChatGPT-based test case generation based on a given Java focal method.
@@
While ChatGPT has been widely used for test case generation, existing work only focuses on taking code as input—not the natural language descriptions as input. Therefore, existing test case generation approaches do not support TDD, as TDD mandates writing test cases based on textual descriptions before writing the actual code. In addition, a systematic literature review by Hou et al. (2023) also confirmed that test case generation approaches to support TDD (i.e., LLM-based Text-to-Testcase generation) remain largely unexplored. To address this research gap, we first conduct a preliminary analysis to empirically investigate the performance of ChatGPT for Text-to-Testcase generation.
@@
Figure 1: An example of test cases generated by the OpenAI’s GPT3.5 model based on a given description input. Syntax correctness is measured by an AST parser, while the test executability is measured by running the test cases against the ground truth code.
@@
An Illustrative Example. We use ChatGPT through the OpenAI API by selecting the recommended GPT-3.5-turbo model endpoint. [^1] The experiment was run in June 2023. margin: Following Chen et al. (2022), our prompt includes the context (i.e., external information or additional context that can steer the model to better responses) and the instruction (i.e., a specific task or instruction that the model should perform). Figure 1 (left) presents an example of the input format (i.e., prompt format) where the context refers to the signature of the Python function with a natural language description in the docstring, without the code implementation (Figure 1 (top right)), while the instruction refers to a natural language comment “# check the correctness of the ‘…’ function” followed by an incompleted assert statement. We randomly select one programming task (Problem #4479) from the testing set of the APPS dataset (Hendrycks et al., 2021) as an illustrative example. Note that a comprehensive evaluation of ChatGPT is provided in Section 5 where we found that our PyTester performs better. For the given Problem #4479, we reformat the data structure into a triplet of $<$function signature + description text, code implementation, test cases$>$. Then, the function signature + description text is input into ChatGPT to generate test cases, while code implementation is only used to evaluate if the generated test cases can be run against the ground truth code.
@@
Observation-1: ChatGPT can generate syntactically correct test cases, yet some of the generated test cases cannot be executed (i.e., failed test cases). In reality, when using ChatGPT to generate test cases based on a given description, it becomes challenging to verify if the generated test cases are correct and aligned with the given description or not. To address this challenge, assuming that the ground truth code is the correct implementation, we execute the test cases against the ground truth code to verify if the ChatGPT-generated test cases are executable or not. Figure 1 (bottom right) shows that four test cases generated by ChatGPT are syntactically correct, but the fourth test case is not executable. In particular, the fourth test case is generated with the test input of ‘([-2,9,-8,8,7], 3)’. Thus, the expected output should be ‘30’, but incorrectly generated as ‘32’. Therefore, this test case will raise an AssertionError. Such syntactically correct, yet non-executable test cases indicate that the test case is still not aligned with the given descriptions. This observation demonstrates that ChatGPT is able to generate syntactically correct test cases, yet fails to generate executable ones (i.e., failed test cases).
@@
Observation-2: When ChatGPT generates non-executable test cases, it becomes challenging to evaluate the completeness and effectiveness of the test cases. Intuitively, effective test cases should exercise all lines of code in the program (i.e., achieving high code coverage), and be able to discover bugs/defects when the program is slightly changed (i.e., the same set of test cases is able to catch the introduced bugs). Code coverage is widely used to measure test completeness  (Pacheco and Ernst, 2007; Fraser and Arcuri, 2011; Schäfer et al., ; Siddiq et al., 2023; Tufano et al., 2020; Alagarsamy et al., 2023; Tang et al., 2023). Thus, a higher code coverage percentage often increases the confidence that the software behaves as intended. However, studies (Hemmati, 2015; Cai and Lyu, 2005) argue that code coverage only measures the extent to which the test cases exercise the lines of code in the program, but not the effectiveness of the test cases. Therefore, mutation testing is proposed to evaluate the effectiveness of the test cases (Cai and Lyu, 2005; Madeyski, 2010; Jia and Harman, 2010; Papadakis et al., 2019). Mutation testing aims to introduce intentional faults (mutations) into the code and check if the existing test cases can detect these faults (mutants)—referring to the fault detection capability of the test cases. Thus, the mutation score is measured as a percentage based on the number of mutants that are killed (i.e. detected by the test cases) versus the total number of mutants introduced. Nonetheless, despite the benefits of code coverage and the extensive research on mutation testing in the literature, another crucial factor is the test case executability. When ChatGPT generates non-executable test cases, such critical metrics (e.g., code coverage and mutation score) are unable to be assessed. Therefore, the models should consider these domain knowledge metrics during training – which is largely overlooked in the existing ChatGPT-based approaches. For instance, although (Chen et al., 2022) proposed a method to re-rank generated code and test cases based on execution feedback, it does not take into account the test case characteristic.
@@
Summary. TDD offers numerous benefits to software development, yet is time-consuming and expensive. Existing test case generation approaches do not yet support TDD practices (i.e., code, not textual description, is treated as input). Our motivating analysis shows that ChatGPT, one of the most powerful language models, still generates non-executable, incomplete, and ineffective test cases, hindering the adoption of ChatGPT and other LLMs for Text-to-Testcase generation in practice. To support TDD practices, this observation highlights the need for a novel automated test case generation approach that can generate syntactically correct, executable, complete, and effective test cases.
3 PyTester